OpenStudy (anonymous):

x = -0.3333333333 + -1.666666667y

OpenStudy (anonymous):

what do you have to do with this function?

OpenStudy (anonymous):

3x+5y=-1

OpenStudy (anonymous):

??

OpenStudy (anonymous):

i have to simplify

OpenStudy (anonymous):

What is the solution of this linear system? 3x + 5y = -1 2x − 5y = 16

OpenStudy (austinl):

Where are we going with this? You gave two different equations. And one of which is a MASSIVE approximation.

OpenStudy (anonymous):

never mind about the 1st ones i just needed help with the one i typed

OpenStudy (austinl):

Which one? There are three different equations. Two if we disregard the first one. The system of equations one?

OpenStudy (wolf1728):

So solve for these two? 3x + 5y = -1 2x − 5y = 16

OpenStudy (anonymous):

guys don't help her, she just wants answers, that is not how this site works

OpenStudy (austinl):

You think I don't realize? It is just that there is so much crap going on here that I am very very confused. I think I am gonna go now.